# Make a Donut: Hierarchical EMD-Space Planning for Zero-Shot Deformable Manipulation with Tools

### v2

@@ Abstract
0
@@ I Introduction
Manipulation of deformable objects remains one of the most intricate challenges in robotics due to the inherent complexity and unpredictable behavior of such objects. Deformable objects can be broadly categorized into two major categories: thin-shell surfaces, such as clothes [1, 2] and ropes [3]; and volumetric objects, such as dough [4, 5]. In this paper, we focus on the latter and study dough manipulation given a set of candidate tools like rolling pin, knife, etc.
Existing works on dough-like volumetric deformable objects majorly rely on a learned dynamic model of the underlying particles  [6, 7, 8]. However, these methods all require a substantial amount of collected or semi-auto-generated demonstrations for training the dynamic models, which poses two critical issues: firstly, the difficulty of obtaining a comprehensive set of demonstrations, particularly for long-horizon tasks; and more importantly, the limited capability of generalizing beyond the scope of the provided demonstrations.
@@
In this study, we delve into the manipulation of dough, a quintessential example of deformable object manipulation [4, 9, 5] with tools. As illustrated in Figure 1, our approach takes a natural language user prompt as input and leverages a large language model (LLM) to formulate a high-level plan detailing the selection of tools and the representation of intermediate subgoals at each phase. While LLMs might not produce precise low-level actions for each timestep, they exhibit proficiency in breaking down intricate tasks into manageable stages. Each of these stages exclusively involves a single tool and a piece of dough.
@@
Through this hierarchical strategy, our system can adeptly tackle novel, intricate tasks devoid of any demonstrations or prior training, given a set of candidate tools. Experimental results demonstrate that our methodology markedly enhances the efficacy of both single-tool and multiple-tool dough manipulation with various tools and can potentially transfer to real-world robotic applications.
@@ III-A Multiple Tool Selection and Hierarchical Planning
To address the challenge of coordinating between different tools in long-horizon tasks, we turn to the capabilities of large language models (LLMs), especially ChatGPT-4 [25]. Our observations indicate that while LLMs may not precisely produce low-level actions, they excel in deconstructing intricate long-horizon tasks into several stages, each centered around a single tool. What’s more, for each of these stages, the LLM can both identify the appropriate tool and generate the corresponding Python code to produce intermediate subgoal point clouds. Presented in the form of particles, these subgoals readily align with the target requirements of our proposed single-tool EMD space planning algorithm.
To guide this process, we devised a prompt template that imparts to the LLM foundational information about available tools and their potential interactions with the dough. Additionally, we introduce a set of guidelines designed to refine and direct the LLM when completing more complex, long-horizon deformable tasks. One important guideline is to force the LLM to give volume-preserving input and output at each stage, so the target is more physically realistic. In addition, we leverage the chain reasoning technique  [26] to help the LLM better deduce the shape parameters to satisfy this constraint. In Table I, we provide the average relative volume change between the LLM’s generated final output and the input dough for all the three evaluated tasks. Details of the prompt template can be found on our project website.
@@
  • The Python code to generate target point clouds. Do remember to add their absolute world location when generating complex shapes.
  • The variable names for the input and output.
  • The location of each piece in a dictionary format with a variable name as the key.
  • The volume of each piece is also in a dictionary format with a variable name as the key.
Building on this, we only extract the generated Python code corresponding to each stage, from which we produce the intermediate subgoal point cloud. Other outputs, though not used, are part of chain reasoning and, therefore contribute to the final quality of the generated subgoal. The subgoal point cloud is then fed directly into our single-tool planning module. Consequently, we are equipped to tackle intricate tasks incrementally, stage by stage, eliminating the need for demonstrations.
@@ III-B Single Tool Planning
Given the current observation and the goal, we pinpoint the subsequent reachable point cloud by executing multiple small iterations (specifically 20 in our experiments) of gradient descent within the Earth Mover Distance (EMD) space. Formally, each iteration is:
@@
Figure 2: EMD-space planning with DiffPhysics-P2P. We find the next reachable target by running small steps within the EMD space. The induced point-to-point correspondence can provide better gradients when optimizing actions through differential physics.
In other words, given the current point cloud observation, denoted as ${\mathbf{p}_{i}}$, where $i$ is the point index, and the subgoal ${\bar{\mathbf{p}}_{i}}$, our objective is to discern the ensuing reachable candidate. This is achieved by incrementally transitioning the current point cloud towards the target within the EMD space. The candidate serves as our model’s prediction of the underlying particle dynamics.
@@
$L=\sum_{i}\|\mathbf{p}_{i}^{\prime}-\mathbf{p}_{i}\|_{1}.$ (2)
@@
The aforementioned EMD planning algorithm demonstrates efficacy when the initial tool position is in proximity to the dough. However, challenges arise when the tool is situated at a considerable distance from the dough, resulting in the algorithm’s inability to find the global minima. To figure out a good initial position for the tool, we leverage the strategy initially proposed in [27]. Specifically, with the present deformation field deduced from the induced point-to-point correspondence (represented as $\mathbf{p}_{i}^{\prime}-\mathbf{p}_{i}$), we employ the following equation to identify the candidate tool position:
$\mathbf{x}^{*}=\arg\max_{\mathbf{x}}\sum_{i}\frac{\|\mathbf{p}_{i}^{\prime}-\mathbf{p}_{i}\|_{1}}{\mathrm{sdf}_{\mathbf{x}}(\mathbf{p}_{i})+\delta}.$ (3)
The numerator encapsulates the point-to-point correspondence loss of the extant point cloud. In contrast, the denominator represents the signed distance field (SDF) of the tool (plus a small delta $\delta$) when positioned at $\mathbf{x}$, evaluated at particle $\mathbf{p}_{i}$, representing distance from the tool to the dough.
@@
Figure 3: Illustration on how tool reset works. By resetting the tool position when no improvement can be made, we can jump out of the local minima and get a better global solution.
Integrating the previously described techniques allows us to first select an optimal initial tool position, followed by iteratively progressing towards the target as per Equation 1. However, even with an advantageous starting tool position, challenges may arise due to the inherent intricacies of differentiable physics. As illustrated in Figure 3, consider a scenario where the task is to use a rolling pin to flatten the dough into a sphere. While initiating from a favorable position, iterating and optimizing candidates in the EMD space could land us at a local minimum. This could result in an uneven texture on one side of the dough, manifesting as a lump. To circumvent this predicament, we reset the tool’s position if no advancement is observed, ensuring an escape from potential local minima. The comprehensive single-tool closed loop planning process is detailed in Algorithm 1.
@@
7:         $\{\mathbf{p}_{i}^{\prime}\}:=\{\mathbf{p}_{i}^{\prime}\}-\alpha\cdot\nabla_{\{\mathbf{p}_{i}^{\prime}\}}\mathrm{emd}(\{\mathbf{p}_{i}^{\prime}\},\{\bar{\mathbf{p}}_{i}\})$triangleright Move particles along the gradient within $\mathrm{emd}$ space
@@
11:         $\mathbf{x}^{*}=\operatorname*{arg\,max}_{\mathbf{x}}\frac{\|\mathbf{p}_{i}^{\prime}-\mathbf{p}_{i}\|_{1}}{\mathrm{sdf}_{\mathbf{x}}(\mathbf{p}_{i})+\delta}$triangleright Find the optimal initial tool position
@@
15:     $\mathbf{a}_{t:t+L}:=\mathbf{0}$triangleright Initialize actions for horizon $L$
@@
17:         $\mathbf{a}_{t:t+L}:=\mathbf{a}_{t:t+L}-\nabla_{\mathbf{a}}\mathrm{Sim}(\sum_{i}\|\mathbf{p}_{i}^{\prime}-\mathbf{p}_{i}\|_{1})$triangleright Differential physics with point-to-point correspondence
@@
19:     Execute $\mathbf{a}_{t:t+L}$ and get the new particle observation $\{\tilde{\mathbf{p}}_{i}\}$triangleright $\{\tilde{\mathbf{p}}_{i}\}$ may differ from $\{\mathbf{p}_{i}^{\prime}\}$
20:     $\mathrm{emd}_{curr}:=\mathrm{emd}(\{\tilde{\mathbf{p}}_{i}\},\{\bar{\mathbf{p}}_{i}\})$triangleright Calculate the actual $\mathrm{emd}$ after executing $\mathbf{a}_{t:t+L}$
21:     if $\mathrm{emd}_{curr}>\mathrm{emd}_{last}$ then $\triangleright$ No progress
22:         $need\_reset=1$triangleright Find another initial position to escape local minimum
@@ III-C Implementation Details on Simulated Experiments
We use a parameter $max\_steps=200$, $K=20$ and $L=40$. We make use of the Adam optimizer with a learning rate set at 0.02 during the differentiable physics computations. For the $\mathrm{emd}$ loss computation, our approach leans on the GeomLoss [28] library available in Pytorch. When searching for the optimal tool position, we consider the tightest axis-aligned bounding box plus a small offset above the dough depending on each tool’s shape.
@@ IV Experiment
In our experiments, we validate the efficacy of our method across four distinct dimensions. In Section IV-A, we deploy our hierarchical LLM-guided planning algorithm to more intricate and unseen long-horizon tasks. In Section IV-B, we conduct simpler tasks involving only one tool to authenticate the effectiveness of our single-tool planning algorithm. In Section IV-C, we perform distinct ablation studies for both single-tool and multiple-tool planning algorithms. In Section IV-D, we translate our simulated actions to a real-world robot to manipulate the actual dough, illustrating the practical applicability of our method from sim to real.
@@
We adopt the metrics in [9, 4] and report the normalized decrease (score) in the Earth Mover Distance (EMD), which is computed using the Sinkhorn divergence as $\frac{\mathrm{emd}(\{\mathbf{p}_{0}\},\{\mathbf{p}^{*}\})-\mathrm{emd}(\{\mathbf{p}_{T}\},\{\mathbf{p}^{*}\})}{\mathrm{emd}(\{\mathbf{p}_{0}\},\{\mathbf{p}^{*}\})}$, where $\{\mathbf{p}_{0}\}$ is the initial point cloud, $\{\mathbf{p}_{T}\}$ is the final point cloud after execution, and $\{\mathbf{p}^{*}\}$ is the ground-truth target 3D point cloud. Additionally, we also calculate the success rates based on pre-set thresholds (Appendix A-A).
@@ Results
The key distinction of our approach lies in its zero-shot learning ability, which enables it to adapt to novel tasks without task-specific fine-tuning or training. This is a significant leap over the sampled-based methods, which may not provide a feasible path or require extensive data for complex shapes. The Large Language Model (LLM) plays a crucial role in our framework by charting a high-level planning path, which serves as a guide for the subsequent execution by the low-level Earth Mover’s Distance (EMD) space planning algorithm. This hierarchical structure is pivotal; the LLM alone cannot translate its generated plans into the raw actions required for robotic manipulation. Conversely, without the strategic direction provided by the LLM, the EMD space planning lacks a coherent objective, struggling to discern what end states are physically plausible for the robot to achieve. PASTA, while effective within its demonstrated scope, requires a dataset to train on in order to sample feasible intermediate states and is thus inherently limited in its ability to generalize to new shapes, e.g., donut. All their modules like VAE, cost predictor, etc., are tailored to their collected training data. This data-driven dependency hinders its application to the more complex tasks our framework successfully tackles.
@@ IV-C Ablation Studies
Figure 7 presents our ablation studies on the removal of the DiffPhysics-P2P, the initial position selection component, or the tool resetting module within the single-tool planning algorithm.
@@
IV-D Real-Robot Experiments
@@
The application of this algorithm to a real-world robot is of immense interest. For this purpose, an experimental setup is established using the UFACTORY xArm 6 and some clay. Given the experimental setting and observations, the proposed planning algorithm is used to generate a trajectory in the simulator, and subsequently, the controller is employed to execute this trajectory in the real world. Unlike that in simulation, we do not have the ground-truth volumetric particle state in real-world experiments. In order to estimate the particle from observations, we setup four cameras from the four orthogonal directions and utilize truncated signed distance function (TSDF) fusion [30] to reconstruct volumetric particles from multi-view depth cameras. In Figure 8, qualitative real-robot trajectories are provided for making a donut. A real-world manipulation demo video is given in the supplementary.

### v3

@@ Abstract
Index Terms:
Task and Motion Planning, Manipulation Planning, Perception for Grasping and Manipulation, AI-Enabled Robotics, Vision Language Model Application
@@ I Introduction
Manipulation of deformable objects remains one of the most intricate challenges in robotics due to the inherent complexity and unpredictable behavior of such objects. Deformable objects can be categorized into two major categories: thin-shell surfaces, like clothes [1, 2] and ropes [3]; and volumetric objects, like dough [4, 5]. In this paper, we focus on the latter and study dough manipulation given a set of candidate tools like rolling pin, knife, etc.
Existing works on dough-like volumetric deformable objects majorly rely on a learned dynamic model of the underlying particles  [6, 7, 8]. However, these methods all require a substantial amount of collected or semi-auto-generated demonstrations for training the dynamic models, which poses two critical issues: firstly, the difficulty of obtaining a comprehensive set of demonstrations, particularly for long-horizon tasks; and more importantly, the limited capability of generalizing beyond the provided demonstrations.
@@
In this study, we delve into the manipulation of dough, a quintessential example of deformable object manipulation [4, 9, 5] with tools. As illustrated in Figure 1, our approach takes a natural language user prompt as input and leverages a LLM to formulate a high-level plan detailing the selection of tools and the representation of intermediate subgoals at each phase. While LLMs might not produce precise low-level actions for each timestep, they exhibit proficiency in breaking down intricate tasks into manageable stages. Each of these stages exclusively involves a single tool and a piece of dough.
@@
Through this hierarchical strategy, our system can adeptly tackle novel, intricate tasks devoid of any demonstrations or prior training, given a set of candidate tools. Experimental results demonstrate that our methodology markedly enhances the efficacy of both single-tool and multiple-tool dough manipulation with various tools and can potentially transfer to real-world robotic applications. To summarize, we propose a novel demonstration-free hierarchical planning method that tackles complex long-horizon deformable manipulation tasks without training. Our contributions are two-fold:
  • We leverage LLMs to manipulate dough by decomposing tasks into distinct stages with achievable subgoals.
  • We employ Differentiable Physics with Point-to-Point correspondence to execute dough manipulation iteratively, eliminating the need for demonstrations.
@@ III-A Multiple Tool Selection and Hierarchical Planning
To address the challenge of coordinating between different tools in long-horizon tasks, we turn to the capabilities of LLMs especially ChatGPT-4 [25]. Our observations indicate that while LLMs may not precisely produce low-level actions, they excel in deconstructing intricate long-horizon tasks into several stages, each centered around a single tool. What’s more, for each of these stages, the LLM can both identify the appropriate tool and generate the corresponding Python code to produce intermediate subgoal point clouds. Presented in the form of particles, these subgoals readily align with the target requirements of our proposed single-tool EMD space planning algorithm (Section III-B).
To guide this process, we devised a prompt template that imparts to the LLM foundational information about available tools and their potential interactions with the dough, which is shared by all tasks. In this work, we consider four different tools: rolling pin, gripper, knife and pole. We require that only one tool be used per stage, which helps ensure that the LLM generates feasible targets matched to the chosen tool. Additionally, we introduce a set of guidelines designed to refine and direct the LLM when completing more complex, long-horizon deformable tasks. One important guideline is to force the LLM to give volume-preserving input and output at each stage, so the target is more physically realistic. In addition, we leverage the chain reasoning technique  [26] to help the LLM better deduce the shape parameters to satisfy this constraint. In Table I, we provide the average relative volume change between the LLM’s generated final output and the input dough for all three evaluated tasks. Importantly, we do not encode any task-specific or object-related information in our prompt, but instead rely on the LLM’s grounding knowledge to select appropriate tools and decompose tasks. We can execute novel tasks by simply adding one line after our prompt. For instance, to make a donut, we only need to concatenate: “the dough is initially a round ball at (0, 0, 0) with radius 1. Make a simple donut.” Detailed prompt can be found on our project website.
@@
  • The Python code to generate target point cloud.
  • The variable name for the generated point cloud.
  • The location of each piece.
  • The volume of each piece.
Building on this, we only extract the generated Python code corresponding to each stage, where the LLM utilizes the numpy library to construct and save the subgoal point cloud into a variable, from which we produce the intermediate subgoal point cloud. Other outputs, though not used, are part of chain reasoning and, therefore, contribute to the final quality of the generated subgoal. The subgoal point cloud is then fed directly into our single-tool planning module. Consequently, we are able to tackle intricate tasks incrementally, stage by stage, without any demonstrations.
@@ III-B Single Tool Planning
Given the current observation and the goal, we pinpoint the subsequent reachable point cloud by executing multiple small iterations (specifically 20 in our experiments) of gradient descent with step size $\alpha$ in the Earth Mover Distance space. Formally, each iteration is:
@@
Figure 2: EMD-space planning with DiffPhysics-P2P. We find the next reachable target by running small steps within the EMD space. The induced point-to-point correspondence can provide better gradients when optimizing actions.
EMD, also known as norm-1 Wasserstein Distance, measures the distribution difference between two point clouds. In our implementation, we compute it with Sinkhorn divergence approximation using GeomLoss [27] Python library. Given the current point cloud observation, denoted as ${\mathbf{p}_{i}}$, where $i$ is the point index, and the subgoal ${\bar{\mathbf{p}}_{i}}$, our objective is to discern the ensuing reachable candidate. This is achieved by incrementally transitioning the current point cloud towards the target within the EMD space. The candidate serves as our model’s prediction of the underlying particle dynamics.
@@
$\mathcal{L}=\sum_{i}\|\mathbf{p}_{i}^{\prime}-\mathbf{p}_{i}\|_{1}.$ (2)
@@
At the beginning of each stage, we use the strategy similar to that in [28] to figure out a good initial position for the tool. Specifically, with the induced point-to-point correspondence (represented as $\mathbf{p}_{i}^{\prime}-\mathbf{p}_{i}$), we find the best tool position $\mathbf{x}^{*}$ by:
$\mathbf{x}^{*}=\arg\max_{\mathbf{x}}(\sum_{i}\frac{\|\mathbf{p}_{i}^{\prime}-\mathbf{p}_{i}\|_{1}}{d(\mathbf{p}_{i},\mathbf{x})+\delta}).$ (3)
The denominator computes the signed distance between point $\mathbf{p}_{i}$ to the closest point on the tool placed at $\mathbf{x}$ (plus small $\delta$ to avoid division by zero). Intuitively, this summation requires positioning the tool at the location where significant deformation is about to happen.
@@
Figure 3: Illustration on how tool reset works. By resetting the tool position when no improvement can be made, we can jump out of the local minima.
As illustrated in Figure 3, consider a scenario where the task is to use a rolling pin to flatten the dough. Sometimes our planning algorithm will get stuck without further progress, leaving a lump. We find that resetting the tool’s position when there is no advancement in the EMD loss helps escape from potential local minima.
The comprehensive single-tool closed loop planning process is detailed in Algorithm 1.
@@
7:         $\{\mathbf{p}_{i}^{\prime}\}:=\{\mathbf{p}_{i}^{\prime}\}-\alpha\cdot\nabla_{\{\mathbf{p}_{i}^{\prime}\}}\mathrm{emd}(\{\mathbf{p}_{i}^{\prime}\},\{\bar{\mathbf{p}}_{i}\})$
@@
11:         $\mathbf{x}^{*}=\operatorname*{arg\,max}_{\mathbf{x}}\frac{\|\mathbf{p}_{i}^{\prime}-\mathbf{p}_{i}\|_{1}}{\mathrm{sdf}_{\mathbf{x}}(\mathbf{p}_{i})+\delta}$
@@
15:     $\mathbf{a}_{t:t+L}:=\mathbf{0}$
@@
17:         $\mathbf{a}_{t:t+L}:=\mathbf{a}_{t:t+L}-\nabla_{\mathbf{a}}\mathrm{Sim}(\sum_{i}\|\mathbf{p}_{i}^{\prime}-\mathbf{p}_{i}\|_{1})$
@@
19:     Execute $\mathbf{a}_{t:t+L}$ and get new particle observation $\{\tilde{\mathbf{p}}_{i}\}$
20:     $\mathrm{emd}_{curr}:=\mathrm{emd}(\{\tilde{\mathbf{p}}_{i}\},\{\bar{\mathbf{p}}_{i}\})$
21:     if $\mathrm{emd}_{curr}>\mathrm{emd}_{last}$ then
22:         $need\_reset=1$
@@
From L4 to L6, we determine the next reachable point cloud by moving particles along the gradient in the EMD space for K iterations. At L9, we compute the optimal initial tool position. Between L13 and L16, we optimize a sequence of actions with a horizon L by running differentiable physics and backpropagating gradients to the actions. Subsequently, we execute the actions and compute the EMD to measure the deviation from the target at L18. If no progress is made (L19), we reset the tool position (L20). At L23, if the system is sufficiently close to the target, we break; otherwise, we continue until the maximum number of timesteps (L2). The model predictive control (MPC) lies in that our model first predicts $\mathbf{p}_{i}^{\prime}$ as the next reachable candidate (L7 of Algorithm 1) — this is the model predictive part, and then uses that prediction to optimize the actions (L15 of Algorithm 1) — this is the optimize and control part
@@ III-C Implementation Details on Simulated Experiments
We use a parameter $max\_steps=200$, $K=20$ and $L=40$. We make use of the Adam optimizer with a learning rate set at 0.02 during the differentiable physics computations.
@@ IV Experiment
In our experiments, we validate the efficacy of our method across four distinct dimensions. In Section IV-A, we deploy our hierarchical LLM-guided planning algorithm to more intricate and unseen long-horizon tasks. In Section IV-B, we conduct simpler tasks involving only one tool to authenticate the effectiveness of our single-tool planning algorithm. In Section IV-C, we perform distinct ablation studies for both single-tool and multiple-tool planning algorithms. In Section IV-E, we translate our simulated actions to a real-world robot to manipulate the actual dough, illustrating the practical applicability of our method from sim to real.
@@
We adopt the metrics in [9, 4] and report the normalized decrease (score) in EMD, which is computed as $\frac{\mathrm{emd}(\{\mathbf{p}_{0}\},\{\mathbf{p}^{*}\})-\mathrm{emd}(\{\mathbf{p}_{T}\},\{\mathbf{p}^{*}\})}{\mathrm{emd}(\{\mathbf{p}_{0}\},\{\mathbf{p}^{*}\})}$, where $\{\mathbf{p}_{0}\}$ is the initial point cloud, $\{\mathbf{p}_{T}\}$ is the final point cloud after execution, and $\{\mathbf{p}^{*}\}$ is the ground-truth target 3D point cloud. Additionally, we also calculate the success rates based on pre-set thresholds (Appendix A-A).
@@ Results
The key distinction of our approach lies in its zero-shot learning ability, which enables it to adapt to novel tasks without task-specific fine-tuning or training. This is a significant leap over the sampled-based methods, which may not provide a feasible path or require extensive data for complex shapes. The LLM plays a crucial role in our framework by charting a high-level planning path, which serves as a guide for the subsequent execution by the low-level EMD space planning algorithm. This hierarchical structure is pivotal; the LLM alone cannot translate its generated plans into the raw actions required for robotic manipulation. Conversely, without the strategic direction provided by the LLM, the EMD space planning lacks a coherent objective, struggling to discern what end states are physically plausible for the robot to achieve. PASTA, while effective within its demonstrated scope, requires a dataset to train on in order to sample feasible intermediate states and is thus inherently limited in its ability to generalize to new shapes, e.g., donut. All their modules like VAE, cost predictor, etc., are tailored to their collected training data. This data-driven dependency hinders its application to the more complex tasks in this paper.
@@ IV-C Ablation Studies
Figure 7 presents our ablation studies on the removal of the DiffPhysics-P2P, the initial position selection component, or the tool resetting module within the single-tool planning algorithm. Without DiffPhysics-P2P means the system directly calculates the EMD loss between $\mathbf{p}_{i}$ (the initial state for this stage) and $\bar{\mathbf{p}_{i}}$ (the LLM-generated goal for this stage), with no point-to-point correspondence. It then optimizes all actions in this stage using $\mathbf{a}=\mathbf{a}-\nabla_{\mathbf{a}}\mathrm{sim}(\mathrm{emd}(\bar{\mathbf{p}}_{i},\mathbf{p}_{i}))$. Our approach differs by gradually moving through the EMD space using P2P loss $\sum_{i}\|\mathbf{p}_{i}^{\prime}-\mathbf{p}_{i}\|_{2}$.
@@
IV-D Time Analysis
Finding the next reachable candidate (L4-L6 of Algorithm 1) takes only 54ms using modern GPUs. Each iteration of the action gradient step with horizon $L=40$ (L15 of Algorithm 1) takes 0.67s, and we perform $J=20$ iterations (L14-L16 of Algorithm 1). Therefore, the average planning time per action is 0.67 * 20 / 40 = 0.34s, or approximately 3Hz.
IV-E Real-Robot Experiments
@@
We set up real-world experiments using the UFACTORY xArm 6 and some clay. Given the experimental setting and observations, the proposed planning algorithm is used to generate a trajectory in the simulator, and subsequently, the controller is employed to execute this trajectory in the real world. Unlike in simulation, we do not have the ground-truth volumetric particle state in real-world experiments. In order to estimate the particle from observations and ensure the initial states in our simulation to be aligned with the real-world observation for each stage, we set up four cameras from the four orthogonal directions and utilize truncated signed distance function (TSDF) fusion [30] to reconstruct volumetric particles from multi-view depth cameras. During each stage, we measure the current particles in real world as the initial state and use LLM-generated subgoals as the target state in the simulator. Using our DiffPhysics-P2P method in the simulator, we gradually move from the initial to target state. We then transfer these simulated actions directly back to the real world. In Figure 8, qualitative trajectories are provided for making a donut.